# InstructMPC: A Human-LLM-in-the-Loop Framework for Context-Aware Control

### v2

@@ 5.3 Application 2: Building Energy Management
6 Conclusion
We have presented the InstructMPC framework that integrates real-time human instructions into Model Predictive Control (MPC) via a human-LLM-in-the-loop approach. This work demonstrates how leveraging contextual information can significantly improve MPC’s predictive accuracy and performance in complex, real-world applications. By dynamically generating context-aware disturbance predictions and refining them using advanced techniques like Proximal Policy Optimization (PPO) and Direct Preference Optimization (DPO), InstructMPC offers improved adaptability and performance across a wide range of applications. For linear dynamics, we established a robust performance guarantee, proving that IMPC achieves a regret bound of $O(\sqrt{T\log T})$. In future work, we aim to extend these results to nonlinear systems and further explore the integration of richer contextual cues to enhance adaptability and robustness in complex environments.
Appendix A Proof of Theorem 4.1
The following lemma characterizing the quadratic cost gap is useful.
@@ 6 Conclusion
For any $\psi_{t}\in\mathbb{R}^{n}$, if at each time $t\in[T]$, a controller $\pi$ implements a control input
$u_{t}^{\pi}=-(R+B^{\top}PB)^{-1}B^{\top}\Big{(}PAx_{t}+\sum_{\tau=t}^{T-1}\left(F^{\top}\right)^{\tau-t}Pw_{\tau}-\psi_{t}\Big{)},$
then the difference between the optimal cost $J^{\star}$ and the algorithm cost $J(\pi)$ is given by:
$J(\pi)-J^{\star}=\sum_{t=0}^{T-1}\psi_{t}^{\top}H\psi_{t},$
where $H\coloneqq B(R+B^{\top}PB)^{-1}B^{\top}$ and $F\coloneqq A-HPA$.
Now, we bound the terms in Equation 9 separately:
@@
Then, rearranging the terms in Equation 10 and noting that the gradients are bounded by Assumption 3,
$2l_{t-k+1}^{\top}(\theta_{t}-\theta^{\star})\leq\frac{\|\theta_{t}-\theta^{\star}\|^{2}-\|\theta_{t+1}-\theta^{\star}\|^{2}}{\eta_{t}}+\eta_{t}G^{2}.$ (11)
Summing above from $t=0$ to $T-1$, for (a),
22 22 $\sum_{t=0}^{T-1}l_{t-k+1}^{\top}(\theta_{t}-\theta^{\star})$
@@
Furthermore, the second term in Equation 9 can be bounded from above as
[…]
@@
$+$ $D\sum_{t=0}^{T-1}\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t}),$ (15)
where we have used (12) and (13) to derive  (14); the last inequality (15) holds by the loss discrepancy in Definition 1. Finally, applying $\eta_{t}={D}/\left({G\sqrt{2(2k-1)(t+1)}}\right)$ to (15),
@@
$+$ $D\sum_{t=0}^{T-1}\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t})+(k-1)GD$
$\leq$ $2GD\sqrt{(k-\frac{1}{2})T}+D\sum_{t=0}^{T-1}\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t})+(k-1)GD.$ (16)
Inequality (16) in the second statement of Theorem 4.1 is obtained by using $\sum_{t=1}^{T}\frac{1}{\sqrt{t}}\leq 2\sqrt{T}$.
Appendix B Proof of Corollary 4.1
@@
$\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t})$ $=\left\|\frac{\partial(\psi_{t}^{\top}H\psi_{t})}{\partial\theta}-\frac{\partial(\hat{\psi}_{t}^{\top}H\hat{\psi}_{t})}{\partial\theta}\right\|$
@@
We have use the fact that $\frac{\partial\psi_{t}}{\partial\theta}=\frac{\partial\hat{\psi}_{t}}{\partial\theta}$ to get (17). Note that if $\mathcal{T}=T-1$, then $\psi_{t}=\hat{\psi}_{t}$, $\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t})=0$. Thus, it suffices to restrict our analysis to the case when $\mathcal{T}=t+k-1$. It follows that
@@
$\leq L\|P\|\sum_{\tau=t}^{t+k-1}\|F\|^{\tau-t}$
$\leq L\frac{\|P\|}{1-\|F\|}.$ (21)
Finally, combining (18),(20), and (21) we have
$\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t})\leq 2\frac{LW\|P\|^{2}\|H\|\|F\|^{k}}{(1-\|F\|)^{2}}.$ (22)

### v3

@@ V-C Application 2: Building Energy Management
VI Conclusion
We have presented the InstructMPC framework that integrates real-time human instructions into Model Predictive Control (MPC) via a human-LLM-in-the-loop approach. This work demonstrates how leveraging contextual information can significantly improve MPC’s predictive accuracy and performance in complex, real-world applications. By dynamically generating context-aware disturbance predictions and refining them using advanced techniques like Proximal Policy Optimization (PPO) and Direct Preference Optimization (DPO), InstructMPC offers improved adaptability and performance across a wide range of applications. For linear dynamics, we established a robust performance guarantee, proving that IMPC achieves a regret bound of $O(\sqrt{T\log T})$. In future work, we aim to extend these results to nonlinear systems, and plan to conduct more experiments under different key parameters, such as the prediction horizon $k$ and the choice of LLM. We will also deploy the InstructMPC framework on our newly deployed physical energy testbed OpenCEM [25] to assess real-world generalization beyond synthetic workloads.
Appendix A: Proof of Theorem IV.1
The following lemma characterizes the quadratic cost gap.
@@ VI Conclusion
For any $\psi_{t}\in\mathbb{R}^{n}$, if at each time $t\in[T]$, a controller $\pi$ implements a control input $u_{t}^{\pi}=-(R+B^{\top}PB)^{-1}B^{\top}\big{(}PAx_{t}+\sum_{\tau=t}^{T-1}\left(F^{\top}\right)^{\tau-t}Pw_{\tau}-\psi_{t}\big{)},$ then the difference between the optimal cost $J^{\star}$ and the algorithm cost $J(\pi)$ is given by $J(\pi)-J^{\star}=\sum_{t=0}^{T-1}\psi_{t}^{\top}H\psi_{t},$ where $H\coloneqq B(R+B^{\top}PB)^{-1}B^{\top}$ and $F\coloneqq A-HPA$.
Now, we bound the terms in Equation˜9 separately:
@@
Then, rearranging the terms in Equation˜10 and noting that the gradients are bounded by Assumption 3,
$2l_{t-k+1}^{\top}(\theta_{t}-\theta^{\star})\leq\frac{\|\theta_{t}-\theta^{\star}\|^{2}-\|\theta_{t+1}-\theta^{\star}\|^{2}}{\eta_{t}}+\eta_{t}G^{2}.$
Summing above from $t=0$ to T1T-1, for (a),
$2$ $\sum_{t=0}^{T-1}l_{t-k+1}^{\top}(\theta_{t}-\theta^{\star})$
@@
Furthermore, the second term in Equation˜9 can be bounded from above as
$\text{(b)}=$ $\sum_{t=0}^{T-1}(l_{t}-l_{t-k+1})^{\top}(\theta_{t}-\theta^{\star})$
@@
$+$ $D\sum_{t=0}^{T-1}\emph{{LD}}(L_{t},\psi_{t}^{\top}H\psi_{t}),$ (14)
where we have used (11) and (12) to derive  (13); the last inequality (14) holds by the loss discrepancy in Definition˜1. Finally, applying $\eta_{t}={D}/\left({G\sqrt{2(2k-1)(t+1)}}\right)$ to (14),
@@
$+$ $D\sum_{t=0}^{T-1}\emph{{LD}}(L_{t},\psi_{t}^{\top}H\psi_{t})+(k-1)GD$
$\leq$ $2GD\sqrt{(k-\frac{1}{2})T}+D\sum_{t=0}^{T-1}\emph{{LD}}(L_{t},\psi_{t}^{\top}H\psi_{t})+(k-1)GD.$
The inequality above in the second statement of Theorem IV.1 is obtained by using $\sum_{t=1}^{T}\frac{1}{\sqrt{t}}\leq 2\sqrt{T}$.
Appendix B: Proof of Corollary IV.1
@@
$\emph{{LD}}(L_{t},\psi_{t}^{\top}H\psi_{t})$ $=\left\|\frac{\partial(\psi_{t}^{\top}H\psi_{t})}{\partial\theta}-\frac{\partial(\hat{\psi}_{t}^{\top}H\hat{\psi}_{t})}{\partial\theta}\right\|$
@@
We have use the fact that $\frac{\partial\psi_{t}}{\partial\theta}=\frac{\partial\hat{\psi}_{t}}{\partial\theta}$ to get (15). Note that if $\mathcal{T}=T-1$, then $\psi_{t}=\hat{\psi}_{t}$, $\emph{{LD}}(L_{t},\psi_{t}^{\top}H\psi_{t})=0$. Thus, it suffices to restrict our analysis to the case when $\mathcal{T}=t+k-1$. It follows that
@@
$\leq L\|P\|\sum_{\tau=t}^{t+k-1}\|F\|^{\tau-t}\leq L\frac{\|P\|}{1-\|F\|}.$
Finally, combining (16),(18), and (Appendix B: Proof of Corollary IV.1) we have
$\emph{{LD}}(L_{t},\psi_{t}^{\top}H\psi_{t})\leq 2\frac{LW\|P\|^{2}\|H\|\|F\|^{k}}{(1-\|F\|)^{2}}.$ (19)